### v4

@@ 1 Introduction
The paper is organized into three parts. First, we start with a quasi-experimental analysis that reveals how methodological choices fundamentally determine merger evaluation conclusions. Standard DiD with fixed effects suggests that all three mergers reduced prices by 4-8% and expanded capacity by 16-29%. However, these pro-competitive effects vanish when controlling for trends or using the first-difference method. This striking sensitivity of the DiD approach motivates our methodological innovations.
@@
Our estimates reveal a concerning temporal progression. The DL-NW merger exhibits decreased coordination (0.141-0.141), but this change is statistically insignificant. However, subsequent mergers increase coordination, with AA-US demonstrating significant competitive deterioration as its conduct increases by $0.107$ (5% level). This escalating pattern validates the DOJ’s concerns about coordinated effects that led it to block the AA-US merger. [^2] See, United States v. US Airways Group, Inc., No. 1:13-cv-01236-CKK (D.D.C. September 5, 2013). These conduct changes largely offset modest efficiency gains, yielding neutral price effects. However, several of the estimates are imprecise because there are limited treated markets–a common challenge in merger evaluation, where only a handful of markets or units are typically treated.
@@ 3.5 DiD Estimates using First Differences
We find little evidence of stable price reductions across any of the three mergers. The DL-NW and UA-CO mergers exhibit modest price impacts that remain largely neutral over the observation period, while the AA-US merger shows no clear downward trend. These findings suggest that potential efficiency benefits from these mergers do not translate into meaningful price reductions.
@@ 4.1 Demand and Supply
On the supply side, we model marginal cost as $C=c(Q,W,I;\boldsymbol{\theta})+\vartheta$, where $Q$ is quantity supplied, $W$ is exogenous cost-shifters independent of $P$, cost shock $\vartheta$ and demand shock $\varphi$, and $I\in\{0,1\}$ is the merger indicator equaling one post-merger in treated market and zero otherwise, and $\boldsymbol{\theta}$ is the vector of cost parameters. Including the merger variable $I$ in the cost function is a parsimonious way to capture merger-induced efficiency, for instance, if $c(Q,W,1;\boldsymbol{\theta})\leq c(Q,W,0;\boldsymbol{\theta})$.
@@
Mergers can also affect competitive conduct. For instance, sustaining supracompetitive pricing may become easier when fewer firms remain post-merger, as monitoring becomes easier, which in turn facilitates tacit coordination. This aspect of the merger suggests conduct may shift toward monopolistic behavior following a merger. To capture this effect, we allow the conduct to change with the merger, and use the notation $\lambda(I)$, such that $\lambda(I=0)=\lambda_{\text{pre}}\leq\lambda(I=1)=\lambda_{\text{post}}$.
@@ B Proofs of the Identification Results
Next, we show that the change in conduct $\Delta\lambda:=\lambda_{\text{post}}-\lambda_{\text{pre}}$ is identified. To this end, we show that $\Delta\lambda$ takes the same value across all observationally equivalent parameter sets. Consider any two observationally equivalent parameter sets $(\theta_{1},\lambda_{\text{pre}},\lambda_{\text{post}})$ and $(\theta_{1}^{*},\lambda_{\text{pre}}^{*},\lambda_{\text{post}}^{*})$ that generate the same observables $\tilde{\gamma}_{0}$ and $\tilde{\gamma}_{1}$. It is easy to check that for both these sets,
$\Delta\lambda=\lambda_{\text{post}}-\lambda_{\text{pre}}=-\alpha_{1}\times(\tilde{\gamma}_{1}-\tilde{\gamma}_{0})=-\alpha_{1}\times(\tilde{\gamma}_{1}^{*}-\tilde{\gamma}_{0}^{*})=\lambda_{\text{post}}^{*}-\lambda_{\text{pre}}^{*}=\Delta\lambda.$
Therefore, $\Delta\lambda$ takes the same value across all observationally equivalent parameter sets, which proves that $\Delta\lambda$ is identified. ∎
@@
Rearranging Equations (B.25) and (B.27) gives $(\theta_{1}-\theta_{1}^{*})=\frac{\lambda_{\text{pre}}}{\alpha_{1}^{1}}-\frac{\lambda_{\text{pre}}^{*}}{\alpha_{1}^{1}}$ and $(\theta_{1}-\theta_{1}^{*})=\frac{\lambda_{\text{pre}}}{\alpha_{1}^{2}}-\frac{\lambda_{\text{pre}}^{*}}{\alpha_{1}^{2}}$, respectively. Equating the two and simplifying the expressions further gives
$\lambda_{\text{pre}}\times\left(\frac{\alpha_{1}^{2}-\alpha_{1}^{1}}{\alpha_{1}^{1}\alpha_{1}^{2}}\right)=\lambda_{\text{pre}}^{*}\times\left(\frac{\alpha_{1}^{2}-\alpha_{1}^{1}}{\alpha_{1}^{1}\alpha_{1}^{2}}\right).$
As $\alpha_{1}^{1}\neq\alpha_{1}^{2}$, the two terms in the parentheses in the equation above are non-zero, and therefore $\lambda_{\text{pre}}=\lambda_{\text{pre}}^{*}$. Applying the same logic as above to Equations (B.26) and (B.28), it follows that we get $\lambda_{\text{post}}=\lambda_{\text{post}}^{*}$.
Finally, we note that because $\lambda_{\text{pre}}=\lambda_{\text{pre}}^{*}$ and $(\theta_{1}-\theta_{1}^{*})=\frac{\lambda_{\text{pre}}}{\alpha_{1}^{1}}-\frac{\lambda_{\text{pre}}^{*}}{\alpha_{1}^{1}}=0$, as shown above, we get $\theta_{1}=\theta_{1}^{*}$. Now that we have established $\lambda_{\text{pre}}=\lambda_{\text{pre}}^{*}$ and $\lambda_{\text{pre}}=\lambda_{\text{pre}}^{*}$. This proves that the parameters $(\theta_{1},\lambda_{\text{pre}},\lambda_{\text{pre}})$ are identified, and are given by
$\theta_{1}=\frac{\alpha_{1}^{1}\tilde{\gamma}_{0}^{1}-\alpha_{1}^{2}\tilde{\gamma}_{0}^{2}}{\alpha_{1}^{1}-\alpha_{1}^{2}};\quad\!\!\lambda_{\text{pre}}=\min\{0,\max\{\alpha_{1}^{1}(\theta_{1}-\tilde{\gamma}_{0}^{1}),1\}\};\quad\!\!\lambda_{\text{post}}=\min\{0,\max\{\alpha_{1}^{1}(\theta_{1}-\tilde{\gamma}_{1}^{1}),1\}\}.$

### v5

@@ 1 Introduction
The paper is organized into three parts. First, we start with a quasi-experimental analysis that reveals how methodological choices fundamentally determine merger evaluation conclusions. Standard DiD with fixed effects suggests that all three mergers reduced prices by 4-8% and expanded capacity by 16-29%. However, these pro-competitive effects vanish when using the first-difference method. This striking sensitivity of the DiD approach motivates our methodological innovations. [^2] Although our contexts differ, de Chaisemartin and D’Haultfœuille (2020) also suggest that the difference between two-way fixed effects and first-difference estimators can inform about the data-generating process.
@@
Our estimates reveal a concerning temporal progression. The DL-NW merger exhibits decreased coordination (0.141-0.141), but this change is statistically insignificant. However, subsequent mergers increase coordination, with AA-US demonstrating significant competitive deterioration as its conduct increases by $0.107$ (5% level). This escalating pattern validates the DOJ’s concerns about coordinated effects that led it to block the AA-US merger. These conduct changes largely offset modest efficiency gains, yielding neutral price effects. However, several of the estimates are imprecise because there are limited treated markets–a common challenge in merger evaluation, where only a handful of markets or units are typically treated.
@@ 3.5 DiD Estimates using First Differences
We find little evidence of stable price reductions across any of the three mergers. The DL-NW and UA-CO mergers exhibit modest price impacts that remain largely neutral over the observation period, while the AA-US merger shows no clear downward trend. Thus any efficiency benefits from these mergers do not translate into meaningful price reductions.
@@ 4.1 Demand and Supply
On the supply side, we model marginal cost as $C=c(Q,W,I;\boldsymbol{\theta})+\vartheta$, where $Q$ is quantity supplied, $W$ is exogenous cost-shifters independent of $P$, cost shock $\vartheta$ and demand shock $\varphi$, and $I\in\{0,1\}$ is the merger indicator equaling one post-merger in treated market and zero otherwise, and $\boldsymbol{\theta}$ is the vector of cost parameters. Including merger $I$ in the cost function captures merger-induced efficiency, for instance, if $c(Q,W,1;\boldsymbol{\theta})\leq c(Q,W,0;\boldsymbol{\theta})$.
@@
Mergers can also affect competitive conduct. For instance, sustaining supracompetitive pricing may become easier when fewer firms remain post-merger, as monitoring becomes easier, which in turn facilitates tacit coordination. This aspect of the merger suggests conduct may shift toward monopolistic behavior following a merger. To capture this effect, we allow the conduct to change with the merger, such that $\lambda(I=0)=\lambda_{\text{pre}}\leq\lambda(I=1)=\lambda_{\text{post}}$.
@@ B Proofs of the Identification Results
Next, we show that the change in conduct $\Delta\lambda:=\lambda_{\text{post}}-\lambda_{\text{pre}}$ is identified. To this end, we show that $\Delta\lambda$ takes the same value across all observationally equivalent parameter sets. Consider any two observationally equivalent parameter sets $(\theta_{1},\lambda_{\text{pre}},\lambda_{\text{post}})$ and $(\theta_{1}^{*},\lambda_{\text{pre}}^{*},\lambda_{\text{post}}^{*})$ that generate the same observables $\tilde{\gamma}_{0}$ and $\tilde{\gamma}_{1}$. It is easy to check that for both these sets, $\Delta\lambda=\lambda_{\text{post}}-\lambda_{\text{pre}}=-\alpha_{1}\times(\tilde{\gamma}_{1}-\tilde{\gamma}_{0})=-\alpha_{1}\times(\tilde{\gamma}_{1}^{*}-\tilde{\gamma}_{0}^{*})=\lambda_{\text{post}}^{*}-\lambda_{\text{pre}}^{*}=\Delta\lambda.$ Therefore, $\Delta\lambda$ takes the same value across all observationally equivalent parameter sets, which proves that $\Delta\lambda$ is identified. ∎
@@
Rearranging Equations (B.25) and (B.27) gives $(\theta_{1}-\theta_{1}^{*})=\frac{\lambda_{\text{pre}}}{\alpha_{1}^{1}}-\frac{\lambda_{\text{pre}}^{*}}{\alpha_{1}^{1}}$ and $(\theta_{1}-\theta_{1}^{*})=\frac{\lambda_{\text{pre}}}{\alpha_{1}^{2}}-\frac{\lambda_{\text{pre}}^{*}}{\alpha_{1}^{2}}$, respectively. Equating the two and simplifying the expressions further gives $\lambda_{\text{pre}}\times\left(\frac{\alpha_{1}^{2}-\alpha_{1}^{1}}{\alpha_{1}^{1}\alpha_{1}^{2}}\right)=\lambda_{\text{pre}}^{*}\times\left(\frac{\alpha_{1}^{2}-\alpha_{1}^{1}}{\alpha_{1}^{1}\alpha_{1}^{2}}\right).$ As $\alpha_{1}^{1}\neq\alpha_{1}^{2}$, the two terms in the parentheses in the equation above are non-zero, and therefore $\lambda_{\text{pre}}=\lambda_{\text{pre}}^{*}$. Applying the same logic as above to Equations (B.26) and (B.28), it follows that we get $\lambda_{\text{post}}=\lambda_{\text{post}}^{*}$.
Finally, we note that because $\lambda_{\text{pre}}=\lambda_{\text{pre}}^{*}$ and $(\theta_{1}-\theta_{1}^{*})=\frac{\lambda_{\text{pre}}}{\alpha_{1}^{1}}-\frac{\lambda_{\text{pre}}^{*}}{\alpha_{1}^{1}}=0$, as shown above, we get $\theta_{1}=\theta_{1}^{*}$. Now that we have established $\lambda_{\text{pre}}=\lambda_{\text{pre}}^{*}$ and $\lambda_{\text{pre}}=\lambda_{\text{pre}}^{*}$. This proves that the parameters $(\theta_{1},\lambda_{\text{pre}},\lambda_{\text{pre}})$ are identified, and are given by $\theta_{1}=\frac{\alpha_{1}^{1}\tilde{\gamma}_{0}^{1}-\alpha_{1}^{2}\tilde{\gamma}_{0}^{2}}{\alpha_{1}^{1}-\alpha_{1}^{2}};\quad\!\!\lambda_{\text{pre}}=\min\{0,\max\{\alpha_{1}^{1}(\theta_{1}-\tilde{\gamma}_{0}^{1}),1\}\};\quad\!\!\lambda_{\text{post}}=\min\{0,\max\{\alpha_{1}^{1}(\theta_{1}-\tilde{\gamma}_{1}^{1}),1\}\}.$